OpenStudy (elise_a18):

Medal and Fan Not a multiple choice question, test, or Exam question What is the formula for finding the height of a square pyramid who's volume is 8 cubic feet and a base length of 2 ft?

OpenStudy (elise_a18):

@vera_ewing @Voltage @Vocaloid

Vocaloid (vocaloid):

good, so we have V = (1/3)(r^2)(h) we have V and r, and we want to find h

OpenStudy (elise_a18):

yep:)

Vocaloid (vocaloid):

so we re-arrange our formula to get h = 3V/(r^2) now plug in V and r

OpenStudy (elise_a18):

I'm having a heard time isolating h

Vocaloid (vocaloid):

ok, want me to walk you through it?

OpenStudy (elise_a18):

Oh! wait!

Vocaloid (vocaloid):

nope, this formula works for a square pyramid a cone would be (1/3)(pi)r^2h

Vocaloid (vocaloid):

if r = the side length of the base, that is

OpenStudy (elise_a18):

Alright, let's start again. I have 2 formulas for a square pyramid \[\frac{ 1 }{ 6 } (b^3) \] and \[\frac{ 1 }{ 3 } (b) (h)\]

Vocaloid (vocaloid):

hm... the first formula only works if the height is equal to the side length of the base

Vocaloid (vocaloid):

lets use the second formula V = (1/3)bh where b is the area of the base and h is the height

OpenStudy (elise_a18):



OpenStudy (elise_a18):

Okay so the formula would be \[\frac{ 1 }{ 3 } (2) (h)\]

Vocaloid (vocaloid):

oh, ok, if 2 is the area of the base, then yes

Vocaloid (vocaloid):

V = (1/3)(2h) now multiply both sides by to "get rid of" the 1/3

OpenStudy (elise_a18):

base length rather :P

OpenStudy (elise_a18):

@Vocaloid

Vocaloid (vocaloid):

wait, why "cubic feet"? that implies 3 dimensions

Vocaloid (vocaloid):

if the side length is 2, then the base is 4

OpenStudy (elise_a18):

yeah, it's a square pyramid. like in my drawing

OpenStudy (elise_a18):

Ohh right! I forgot that BL= S^2

Vocaloid (vocaloid):

right, so if b = 4 then V = (1/3)(4)(h) know how to isolate h now?

OpenStudy (elise_a18):

Yeah I did it on my white board. Would it be h= \[\frac{ 1 }{ 3 } *8\frac{ 4 }\]

OpenStudy (elise_a18):

oh. O-o

Vocaloid (vocaloid):

nope, not quite

OpenStudy (elise_a18):



Vocaloid (vocaloid):

we want to "get rid of" the 1/3 fraction so we multiply both sides by 3

OpenStudy (elise_a18):

not like this?

Vocaloid (vocaloid):

V = (1/3)(4)(h) 3V = 4h with me so far?

OpenStudy (elise_a18):

yes. So it would be V=8 right? because the vol is already given?

OpenStudy (elise_a18):

OH! it's 6?

Vocaloid (vocaloid):

nope, remember, we want to find h

Vocaloid (vocaloid):

3V = 4h now divide both sides by 4

OpenStudy (elise_a18):

i though we need to divide 3V by 4 to isolate h?

Vocaloid (vocaloid):

right

Vocaloid (vocaloid):

h = 6, right

Vocaloid (vocaloid):

didn't see your answer up there ;P

OpenStudy (elise_a18):

Yeay! Thank you so much! I know I'm kinda slow but I love it when I can actually understand X)

OpenStudy (elise_a18):

i was seeing that V is substituted by 8